# Tuning Large Multimodal Models for Videos using Reinforcement Learning from AI Feedback

### v1

@@ 4.1 Experimental Setup
Model settings.
We initiate training of the VLM-SFT, based on the pre-trained LLaVA image-text large multimodal models Liu et al. (2023a), leveraging a diverse video-text instruction-tune dataset Muhammad Maaz and Khan (2023); Xiao et al. (2021); Li et al. (2020); Voigtlaender et al. (2023). This training incorporates a curriculum approach as outlined in Sec. 3.3. Upon establishing the VLM-SFT, we proceed to train the reward model (RM) using the VLM-SFT for both its foundation and the generation of preference data. Subsequently, we implement reinforcement learning with the trained RM, leading to the development of the VLM-RLAIF. In the reinforcement learning phase, the policy model is initialized from the VLM-SFT, while the value model employs the RM.
@@
For the SFT dataset, we utilize the open-sourced video-text instruction-tune dataset (80k) Muhammad Maaz and Khan (2023); Li et al. (2023a) and video question answering datasets (67k) Xiao et al. (2021); Li et al. (2020). Additionally, we generate object-centric narrative video-text instruction-tune dataset (180k) for training the VLM-SFT, as mentioned in Sec. 3.3. For two-stage curriculum learning, we divide the instruction-tune data for SFT into two groups based on the difficulty; easy (214k) and hard (113k) data. To train the reward model (RM), we first generate responses from existing instruction-tune data Muhammad Maaz and Khan (2023) and generate preferences using them (40k). Then, we once again utilize the existing instruction-tune dataset (100k) Muhammad Maaz and Khan (2023) for reinforcement learning with the trained RM for initialization.
Evaluation details.
We evaluate our VLM-RLAIF on various video benchmarks; zero-shot video questions and video-based generative benchmarks following the prior work Muhammad Maaz and Khan (2023); Lin et al. (2023); Li et al. (2023c) and text-to-video-retrieval and action recognition as proposed in prior work Li et al. (2023b).
@@
Table 2: Qualitative comparison between different VLMMs on zero-shot video question answering benchmark. Our VLM-RLAIF outperforms previous work across three video-question answering benchmarks.
@@
Table 3: Qualitative comparison between different VLMMs on zero-shot text-to-video (T2V) retrieval and action recognition. Following prior work Li et al. (2023b), we assess our proposed VLM-RLAIF on zero-shot T2V and action recognition for comprehensive evaluation of the VLMMs. $\dagger$} denotes reproduced results using the author’s implementation.
Figure 5: Effect of various preference data size on video-based generative benchmark. VLM-RLAIF performance increase with the number of collected preference data.
@@ 4.2 Quantitative Analysis
We quantitatively evaluate our proposed VLM-RLAIF on various video benchmarks following prior work: video-based generative benchmark, zero-shot video question answering Muhammad Maaz and Khan (2023); Lin et al. (2023); Li et al. (2023c), text-to-video retrieval and action recognition Li et al. (2023b).
@@
We evaluate VLMMs on the video-based generative performance benchmark Muhammad Maaz and Khan (2023) that measures five criteria: correctness, detail, context, temporal understanding, and consistency of generated text. In specific, these assess the relevance of the model’s output to the video content, its capacity to capture essential details and contextual information, its understanding of temporal sequences, and the consistency in responding to varied yet related queries. As indicated in Tab. 1, the VLM-RLAIF performs on par with GPT-4V OpenAI et al. (2023) and outperforms previous approaches and the VLM-SFT.
@@
To assess the general capabilities of various video-text Large Language Models (LLMs), we conduct a quantitative evaluation of their video question answering (VideoQA) abilities using three datasets: MSVD-QA, MSRVTT-QA Xu et al. (2017), and ActivityNet-QA Yu et al. (2019). We adopted the evaluation pipeline from prior work Muhammad Maaz and Khan (2023), which utilizes ChatGPT-assisted evaluation, to gauge the zero-shot VideoQA performance. The results, as shown in Table 2, indicate that the VLM-RLAIF significantly outperforms previous models. Notably, VLM-RLAIF exceeds VLM-SFT by 9.2%, 10.6%, and 13.2% in accuracy and by 0.4, 0.4, and 0.3 in score across all datasets.
@@
For this evaluation, we follow the procedure proposed in prior work Li et al. (2023b), which compare CLIP score between generated description and ground-truth caption. Table 3 illustrates the summarized performance comparison to various VLMMs. In two datasets, i.e., MSVD and MSRVTT, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
@@
Following the VLMMs evaluation procedure proposed in prior work Li et al. (2023b), we conduct zero-shot action recognition task using two benchmark datasets, e.g., UCF101 and HMDB51. Table 3 illustrates the summarized performance comparison to various VLMMs. In two datasets, clearly the proposed VLM-RLAIF are much better than other methods including the VLM-SFT.
@@
[A] [B] [C] Corr. Det. Cont. Temp. Cons.
@@
Table 4: In-depth analysis for the VLM-SFT training procedure. ‘[A]’ indicate the multimodal instruction-tune dataset proposed in Muhammad Maaz and Khan (2023); Li et al. (2023a). ‘[B]’ represents the use of a human-labeled video question answering dataset Xiao et al. (2021); Li et al. (2020), while ‘[C]’ refers to the use of an object-centric video narrative instruction-tune dataset (Appendix Sec. B). ‘Curr. learning’ indicates the application of curriculum learning as detailed in Sec. 3.3.
@@
Info. Corr. Det. Cont. Temp. Cons.
@@
Table 5: Effect of context information on video-based generative performance benchmark Muhammad Maaz and Khan (2023). We investigate the efficacy of using context information for reward modeling (Sec. 3.2). ‘Context Info.’ indicates the use of context in preference selection. ‘#\# Clips’ refers to the the number of times we divide the video into segments to generate the context information.
@@ 4.3 Detailed Analysis
For a detailed analysis, we leverage the video-based generative benchmark Muhammad Maaz and Khan (2023) specifically, as it is well suited to assess the wide-ranging capabilities of VLMMs.
@@
We first empirically support the effectiveness of augmenting the SFT dataset with additional instruction-following dataset (Sec. 3.3). The first and second rows of Tab. 4 illustrate the benefits of incorporating this additional dataset in improving video-based generative performance. On top of that, the application of curriculum learning significantly boosts performance, demonstrating the efficacy of employing a curriculum learning strategy during the SFT process, as illustrated in the third row of Tab. 4.
Effect of various preference data size.
Our framework’s strength lies in generating synthetic preference feedback in large quantities. To evaluate this capability, we escalated the amount of preference data from 10k – the same quantity utilized in the multimodal image-text RLHF framework Sun et al. (2023a) – to 40k. Figure 5 shows a steady improvement in performance with the increase in data size.
@@
Table 5 presents a study on the impact of context information on the performance of the video-based generative benchmark discussed in Sec. 3.2. This reveals that the use of context information during preference selection improves model performance, as demonstrated in the second and third rows in Table 5. Moreover, segmenting the video into multiple clips, specifically three, and using their aggregated caption as context information further enhances performance, as indicated in the third and fourth rows. We conjecture that segmenting video into clips and providing detailed captions for each clip enables the model to capture more detailed nuances, thereby enriching the contextual clarity for reward modeling and boosting the VLM-RLAIF’s performance.
Figure 6: Samples of VLM-RLAIF and VLM-SFT in video understanding. Two samples demonstrate superior capability of the VLM-RLAIF. Red color indicates visually un-grounded expressions considering video content and blue color well grounded to the visual content.
@@ 4.4 Qualitative Analysis
VLM-SFTvs. VLM-RLAIF.
Figure 6 qualitatively compares the performance of VLM-SFT and VLM-RLAIF across two video samples, highlighting their multimodal understanding capabilities. VLM-RLAIF consistently yields more accurate answers than VLM-SFT, as seen in the detailed recognition of an athlete’s attire in the first example, marked in red and blue. The second example further affirms VLM-RLAIF’s superiority in generating precisely grounded responses to the visual input, where VLM-SFT falls short. More examples are in the Appendix Fig. 11.
@@ 5 Conclusion
We introduce a novel alignment strategy for VLMMs, termed VLM-RLAIF, utilizing reinforcement learning from AI feedback. To enhance multimodal alignment, we also present a context-aware reward modeling technique, enabling AI to generate feedback for self-improvement. Additionally, we expand the instruction-tune dataset for SFT and adopt a straightforward curriculum training approach. The VLM-RLAIF significantly surpasses previous models across multiple multimodal video-text understanding benchmarks, including video-based generative tasks, zero-shot VideoQA, text-to-video retrieval, and action recognition.
Appendix A Quantitative Comparison between Different SFT Model Sizes
We assess the performance of the VLM-SFT using varying LLM sizes, specifically 7B and 13B, across video question answering (VideoQA) and video-based generative benchmarks, detailed in Table 6. The results indicate similar outcomes in the zero-shot VideoQA tasks across both model sizes. However, in video-based generative tasks, a clear trend emerges: larger LLM sizes lead to improved performance. This suggests that the increased model capacity inherent in larger LLMs enhances the model’s ability to capture and generate more complex and nuanced video content representations. Thus, we adopt the 13B model for the reward model (Sec. 3.2), with empirical evidence provided in Tables 7 and 8.
Appendix B Details About Object-Centric Instruction-Tune Data
@@
Appendix C Input Prompts for Preference Modeling
We leverage three types of input prompts for context-aware preference modeling, as discussed in Sec. 3.2. Figures 89 and 10 show the detailed prompt example for each stage, respectively.
Appendix D Various LLM Size for Reward Model
In Tables 7 and 8, we assess the performance of VLM-RLAIF using varying reward model(RM) sizes, each initialized from VLM-RLAIF-7B and VLM-RLAIF-13B. In both settings, we use 7B VLM-SFT to initialize the policy model. For both sizes of RM, our RLAIF method showed a performance gain with a large margin from VLM-SFT across all evaluation methods. Specifically, RLAIF with 7B RM acquired 5-12% improvement on zero-shot video question answering. Scaling up the RM from 7B to 13B result in more improved performance, except for text-to-video retrieval task R@5 metric on MSR-VTT.
Figure 7: Designed input prompt for object-centric instruction-tune dataset generation from video narrative dataset and the data sample. In (a), the prompt we used to generate instruciton-following dataset is displayed in a white box. It includes the system prompt, a task definition, guidelines, principles, and an example set of input and desired output. Using this prompt with ChatGPT-3.5-turbo, we create an additional instruction following datasets. The resulting example is visualized in (b). The question covers details of the video scenes, such as the number and appearance of parrots and the action of each parrot.

### v3

@@ 4.1 Experimental Setup
Model details.
We initiate training of the VLM-SFT, building on a pre-trained image-text model Liu et al. (2023a), with various video-text instruction-tune datasets. In particular, we employ a video projection layer consisting of two linear layers with ReLU activation function between them. Upon establishing the VLM-SFT, we train the RM using the VLM-SFT for both its foundation and the generation of preference data. Subsequently, we train the RM using reinforcement learning (VLM-RLAIF). The policy model is initialized from the VLM-SFT, while the value model is initialized from the RM.
@@
For the SFT dataset, we utilize the open-source video-text instruction-tune dataset (80k) Muhammad Maaz and Khan (2023); Li et al. (2023b) and video question answering datasets (67k) Xiao et al. (2021); Li et al. (2020). More importantly, we generate object-centric narrative video-text instruction-tune dataset (180k) for training the VLM-SFT (Sec. 3.2). For the two-stage curriculum learning, we divide the instruction-tune data for SFT into two groups based on the difficulty; easy (214k) and hard (113k) data. To train the RM, we first generate responses from existing instruction-tune data Muhammad Maaz and Khan (2023) and generate preferences using them (40k). Then, we again use the existing instruction-tune dataset (100k) Muhammad Maaz and Khan (2023) for RL with the trained RM.
@@
Table 2: Quantitative comparison between different VLMMs on zero-shot video question answering benchmark. VLM-RLAIF outperforms previous work across three video-question answering benchmarks.
@@
Table 3: Quantitative comparison between different VLMMs on zero-shot text-to-video (T2V) retrieval and action recognition. Following Li et al. (2023c), we evaluate our proposed VLM-RLAIF on zero-shot T2V retrieval and action recognition. $\dagger$}: reproduced by the authors’ implementation.
Figure 5: Effect of preference data size on video-based generative benchmark. VLM-RLAIF’s performance improves across five metrics as the amount of collected preference data increases. The metrics evaluate whether the model generates correct descriptions of the video, includes detailed explanations, remains contextual, demonstrates temporal understanding, and provides consistent responses to variations of the same question.
@@ 4.2 Quantitative Analysis
We evaluate our proposed VLM-RLAIF on various video benchmarks including video-based generative benchmark, zero-shot video question answering Muhammad Maaz and Khan (2023); Lin et al. (2023); Li et al. (2023d), text-to-video retrieval, and action recognition Li et al. (2023c).
@@
We evaluate VLMMs on the video-based generative performance benchmark Muhammad Maaz and Khan (2023) that measures five criteria of generated text. In specific, these assess the relevance of the model’s output to the video content, its capacity to capture essential details and contextual information, its understanding of temporal sequences, and the consistency in responding to varied yet related queries. As shown in Tab. 1, the VLM-RLAIF performs on par with GPT-4V OpenAI et al. (2023), which requires much more computational resources than ours (i.e., not a fair comparison), and outperforms previous approaches and the VLM-SFT.
@@
To evaluate the reasoning ability of VLLMs, we conduct a quantitative evaluation of video question answering (VideoQA) abilities on three datasets Xu et al. (2017); Yu et al. (2019), following Muhammad Maaz and Khan (2023).
The results, as shown in Table 2, indicate that the VLM-RLAIF significantly outperforms previous approaches, including VLM-SFT. Notably, VLM-RLAIF exceeds VLM-SFT by 9.2%, 10.6%, and 13.2% in accuracy and by 0.4, 0.4, and 0.3 in score across all datasets. We believe that the better visually-aligned response generated from the VLM-RLAIF improves the performance (see quantitative analysis in Figures 6 and 11).
@@
For this task, we follow the procedure proposed in Li et al. (2023c), which compares CLIP score between generated description and ground-truth caption. Table 3 illustrates the summarized performance comparison to various VLMMs. In the two datasets, i.e., MSVD and MSRVTT, the proposed VLM-RLAIF clearly outperforms other methods including our VLM-SFT by the help of better alignment by the proposed components.
@@
Following the VLMMs evaluation procedure proposed in Li et al. (2023c), we conduct zero-shot action recognition task using two benchmark datasets, e.g., UCF101 and HMDB51. We summarize results of various VLMMs in Tab. 3. In two datasets, the proposed VLM-RLAIF again clearly outperforms other methods including our VLM-SFT.
@@
[A] [B] [C] Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
@@
Table 4: In-depth analysis for the VLM-SFT training procedure. ‘[A]’ indicate the multimodal instruction-tune dataset proposed in Muhammad Maaz and Khan (2023); Li et al. (2023b). ‘[B]’ represents the use of a human-labeled video question answering dataset Xiao et al. (2021); Li et al. (2020), while ‘[C]’ refers to the use of an object-centric video narrative instruction-tune dataset (Appendix Sec. A). ‘Curr. learning’ indicates the curriculum learning (Sec. 3.2).
@@
Info. Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
@@
Table 5: Effect of context information on video-based generative performance benchmark. We investigate the efficacy of using context information for reward modeling (Sec. 3.1). ‘Context Info.’ indicates the use of context in preference selection. ‘#\# Clips’ denotes the number of segments into which we divide the video to generate the context information.
Methods LLM Size Video-based Generative Perf.
Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
VLM-SFT 7B 2.32 2.53 3.03 2.16 2.23
VLM-SFT 13B 2.64 2.73 3.28 2.38 2.44
Table 6: Quantitative comparison between different sizes of the VLM-SFT. We assess the performance of VLM-SFT with varying LMM sizes, specifically 7B and 13B, on video-based generative benchmarks. We conduct this evaluation without the integration of augmented instruction-tune data and the implementation of two-stage curriculum learning.
Methods LLM Size Video Question Answering Video-based Generative Perf.
Policy Reward MSVD MSRVTT ActivityNet Corr. $\uparrow$ Det. $\uparrow$ Cont. $\uparrow$ Temp. $\uparrow$ Cons. $\uparrow$
Model Model Acc. Score Acc. Score Acc. Score
LLaMA-VID (7B) - - 69.7 3.7 57.7 3.2 47.4 3.3 2.96 3.00 3.53 2.46 2.51
VLM-SFT 7B - 67.2 3.6 52.4 3.0 44.1 3.2 2.79 2.82 3.37 2.28 2.49
VLM-RLAIF 7B 7B 75.1 3.9 61.0 3.3 56.1 3.4 3.47 3.14 3.87 3.05 3.30
VLM-RLAIF 7B 13B 76.4 4.0 63.0 3.4 57.3 3.5 3.63 3.25 4.00 3.23 3.32
Table 7: Quantitative comparison between different size of policy model and reward model for the VLM-RLAIF. We evaluate the VLM-RLAIF with different model size for policy model and reward model on zero-shot video question answering and video-based generative benchmark.
@@ 4.3 Detailed Analysis
For a detailed analysis, we use the video-based generative benchmark Muhammad Maaz and Khan (2023) specifically, as it is well suited to evaluate the wide-ranging capabilities of VLMM, i.e., focusing on response relevance, detail and context capture, temporal understanding, and consistency across queries.
@@
We first empirically support the effectiveness of augmenting the SFT dataset with additional instruction-following dataset (Sec. 3.2). The first and second rows of Tab. 4 illustrate the benefits of incorporating this additional dataset in improving performance. On top of that, the application of curriculum learning significantly improves performance, implying the efficacy of curriculum learning for the SFT process (the third row of Tab. 4).
Effect of preference data size.
Our method’s strength lies in generating synthetic preference feedback in large quantities. To study the benefit of large-sized data, we sweep the size of preference data from 10k – the same quantity utilized in the multimodal image-text RLHF framework Sun et al. (2023a) – to 40k. As expected, we observe monotonic increases in performance by the increase in data size, as shown in Figure 5.
Methods LLM Size T2V Retieval Action Recognition
Policy Reward MSVD MSRVTT UCF101 HMDB51
Model Model R@1 R@5 R@1 R@5 Top-1 Top-5 Top-1 Top-5
LLaMA-VID (7B) - - 27.28 53.40 17.00 35.10 56.58 82.79 38.85 65.27
VLM-SFT 7B - 26.65 54.27 13.10 30.50 53.03 80.34 38.58 62.37
VLM-RLAIF 7B 7B 33.73 61.95 20.80 42.90 61.09 85.15 43.86 65.88
VLM-RLAIF 7B 13B 36.03 63.40 21.00 40.70 62.83 85.86 44.75 68.37
Table 8: Quantitative comparison between different size of policy model and reward model for the VLM-RLAIF. We evaluate the VLM-RLAIF with different model size for policy model and reward model on zero-shot text-to-video retrieval and action recognition tasks.
Figure 6: Samples of VLM-RLAIF and VLM-SFT in video understanding. Two samples demonstrate better capability of the VLM-RLAIF. Red color indicates visually un-grounded expressions considering video content and blue color indicates well grounded expressions to the video.
@@
Table 5 presents a study on the impact of context information. This reveals that the use of context information during preference selection improves performance, as demonstrated in the second and third rows in Tab. 5. Moreover, segmenting the video into multiple clips, specifically three, and using their aggregated captions as context further improves performance (see the third and fourth rows). We conjecture that segmenting video into clips with detailed captions enables the model to capture detailed nuances, thereby enriching the contextual clarity for reward modeling.
Comparison between different SFT model sizes.
Table 6 shows a clear performance margin between VLM-SFT with different LLM sizes, specifically 7B and 13B, in video-based generative tasks. Larger LLM sizes lead to improved performance, suggesting that increased model capacity enhances the ability to capture and generate complex video content. Thus, we adopt the 13B model for the reward model (Sec. 3.1), with supporting evidence in Tables 7 and 8.
Various LLM size for reward model
Tables 7 and 8 show the performance of VLM-RLAIF with different RM sizes, initialized from VLM-RLAIF-7B and VLM-RLAIF-13B. In both cases, the policy model is initialized with VLM-SFT-7B. Our RLAIF method outperforms VLM-SFT significantly across all evaluation methods. Specifically, RLAIF with 7B RM achieves a 5-12% improvement in zero-shot video question answering. Scaling up the RM from 7B to 13B further improves performance, except for the text-to-video retrieval task R@5 metric on MSR-VTT.
@@ 4.4 Qualitative Analysis
We now qualitatively compares the performance of VLM-SFT and VLM-RLAIF, highlighting their multimodal understanding capabilities in Figure 6. VLM-RLAIF consistently yields more accurate answers than VLM-SFT, as shown in the detailed recognition of the attire of an athlete in the first example, marked in red and blue. The second example further affirms VLM-RLAIF’s benefit in generating better grounded responses to the visual input, where VLM-SFT falls short. More examples are in the Appendix Fig. 11 for the space sake.
@@ 5 Conclusion
We propose a novel alignment strategy for VLMMs, termed VLM-RLAIF, that uses reinforcement learning from AI feedback. To improve multimodal alignment, we propose a context-aware reward modeling, enabling AI to generate feedback for self-improvement. In addition, we expand the instruction-tune dataset for SFT and adopt a curriculum-based training approach, that is particularly effective in the gradual learning of complex video-text relationships. In our empirical validations, the VLM-RLAIF significantly outperforms previous models in multiple multimodal video-text understanding benchmarks, which implies good generalization performance across tasks.
Limitations
Given that our approach utilizes feedback synthesized by the AI model, the effectiveness of our proposed VLM-RLAIF largely depends on the quality of the AI model’s generated responses. In light of recent studies exploring the use of artificially generated data Koo et al. (2023); Das et al. (2024), we believe that there needs further research to enhance the quality of synthetically generated data, thereby establishing a more reliable RLAIF system.
In addition, although we have evaluate our model across a range of benchmarks for VLMMs, e.g., videoQA, video-based generative tasks, retrieval, and recognition tasks, there are other tasks such as temporal reasoning Liang et al. (2022) that are necessary for VLMMs to be effectively applied in real-world scenarios. Application of our method to these tasks would be a great future research avenue.
Acknowledgment
This work was partly supported by the NRF grant (No.2022R1A2C4002300, 20%) and IITP grants (No.RS-2022-II220077 (10%), No.RS-2022-II220113 (25%), No.RS-2022-II220959 (10%), No.RS-2022-II220871 (10%), No.RS-2022-II220113 (10%, Yonsei AI), No.RS-2021-II211343 (5%, SNU AI), No.RS-2020-II201361 (5%, AI Innov. Hub), No.RS-2022-11220951 (5%)) funded by the Korea government (MSIT), NCSOFT and Artificial intelligence industrial convergence cluster development project funded by MSIT and Gwangju Metropolitan City.
Appendix
Appendix A Details About Object-Centric Instruction-Tune Data
@@
Appendix B Input Prompts for Reward Modeling
We leverage three types of input prompts for context-aware reward modeling, as discussed in Sec. 3.1. Figure 8 presents a designed input prompt given to our method, resulting in two responses. Furthermore, Figure 9 illustrates the detailed input prompt used to select a preference between two responses. Additionally, Figure 10 displays the input prompt employed for generating a reward score.
Figure 7: Designed input prompt for object-centric instruction-tune dataset generation from video narrative dataset and the data sample. In (a), the prompt we used to generate instruction-following dataset is displayed in a white box. It includes the system prompt, a task definition, guidelines, principles, and an example set of input and desired output. Using this prompt with ChatGPT-3.5-turbo, we create an additional instruction following datasets. The resulting example is visualized in (b). The question covers details of the video scenes, such as the number and appearance of parrots and the action of each parrot.
@@ Appendix B Input Prompts for Reward Modeling
Figure 12: Predictions of SFT model and the predicted reward. We present qualitative examples of questions and two responses generated from the VLM-SFT, and the reward given to each answer using the trained RM. The reward is higher for the answer with more accurately grounded answers, as highlighted in blue for accurate responses and red for less accurate contents.